theyadoreshayyy:

Triangle RST has vertices located at R (2, 3), S (4, 4), and T (5, 0). Part A: Find the length of each side of the triangle. Show your work. Part B: Find the slope of each side of the triangle. Show your work. Part C: Classify the triangle. Explain your reasoning.

new2luv:

do you have a screen shot of the triangle

TyDaGreatest:

Explain what the problem is deeper for me

theyadoreshayyy:

all i need id for somebody to graph the pints on the grid and help me find the slope of the triangle

theyadoreshayyy:

is*

TyDaGreatest:

ok I can help

new2luv:

 so then you need me to put the points on here

theyadoreshayyy:

i need them on a grid

TyDaGreatest:

TyDaGreatest:

that should help

theyadoreshayyy:

how do i find the slope

Woodstock:

Let me formally answer this one, your graph isn't helpful.

TyDaGreatest:

lets see what u got

Woodstock:

1 attachment
Woodstock:

Alright, so when finding slope you;

Woodstock:

You start at (5,0).

Woodstock:

Or at (2,3), whatever you perfer.

theyadoreshayyy:

we can start at 5,0

Woodstock:

If you start at (5,0) you see you go 6 left, and 6 up

Woodstock:

Which gives you a slope of \[\frac{ 6 }{ 6 }=1\] So, your slope is 1.

Woodstock:

I hope that helped, I have to get back to EOC testing.

Woodstock:

See you guys.

theyadoreshayyy:

thank yuuuu

theyadoreshayyy:

waitttt, what about the lengths ? how do i find that

theyadoreshayyy:

@snowflake0531

snowflake0531:

For the lengths, you have to use the distance formula

1 attachment
snowflake0531:

For slope, just do the regular thing, slope between points And for the type of triangle, just graph it on desmos, and what do you think

theyadoreshayyy:

i need help solving the distance formula

snowflake0531:

PLug the points into it, the 4 numbers, square the ones that need to be squared, square root it Looking at R (2, 3), S (4, 4) We can do \[\sqrt ^{(4-2)^2 + (4-3)^2}\]

snowflake0531:

That failed

snowflake0531:

\[\sqrt(4-2)^2+(4-3)^2\] there, but they're all under the square root thing tho o-O

theyadoreshayyy:

so the answer is 3

snowflake0531:

How- It would be \[\sqrt(4+1) = \sqrt5\]

theyadoreshayyy:

snowflake0531:

._. I told you all the stuff is under the square root, when you put it into the calculator, it'll separate bc of the way i wrote it

theyadoreshayyy:

ohhhhh, okay